Problem 24-11 Consider the following information regarding the performance of a money manager in a recent month. The table rea-1. What was the manager’s return in the month? (Do not round intermediate calculations. Input all amounts as positive values. Round your answer to 2 decimal places.)

a-2. What was her overperformance or underperformance? (Do not round intermediate calculations. Input all amounts as positive values. Round your answer to 2 decimal places.)

b. What was the contribution of security selection to relative performance? (Do not round intermediate calculations. Round your answer to 2 decimal places. Negative amount should be indicated by a minus sign.)

c. What was the contribution of asset allocation to relative performance? (Do not round intermediate calculations. Round your answer to 2 decimal places. Negative amount should be indicated by a minus sign.)
Answer #1

a-1) Manager's return = Equity Return*Equity Weight + Bond Return*Bond Weight + Cash Return*Cash Weight

= 2.4%*0.6+1.4%*0.3+0.5%*0.1 = 1.91%

a-2)

Benchmark's Return = Equity Return*Equity Weight + Bond Return*Bond Weight + Cash Return*Cash Weight

= 2.9%*0.7+1.6%*0.2+0.5%*0.1 = 2.4%

Manager's Performance = Manager's return - Benchmark's Return = 1.91%-2.4% = - 0.49% (under performance)

b)

Differential Return (Manager - benchmark) (1) Manager's portfolio weight (2) Contribution to performance (1)*(2)
-0.50% 0.6 -0.30%
-0.20% 0.3 -0.06%
0.00% 0.1 0.00%
-0.36%

Differential Returns: Manager's return in equity - Benchmark Return in Equity = 2.4%-2.9% = - 0..5%

Manager's return in Bonds - Benchmark Return in Bond = 1.4%-1.6% = - 0.2%

Manager's return in cash - Benchmark Return in cash = 0.5%-0.5% =0%

Contribution to performance to security Selection= Differential returns*weight in manager portfolio = -0.5%*0.6 - 0.2%*0.3 + 0%*0.1 = - 0.36%

c)

Differential weight (Manager - benchmark) (1) Index return (2) Contribution to performance of Asset allocation (1)*(2)
-0.1 2.90% -0.29%
0.1 1.60% 0.16%
0 0.50% 0.00%
-0.13%

Differential Returns: Manager's Weight in equity - Benchmark Weight in Equity = 0.6-0.7= - 0.1

Manager's Weight in Bonds - Benchmark Weight in Bond = 0.3-0.2 = 0.1

Manager's Weight in cash - Benchmark Weight in cash = 0.1-0.1 =0

Contribution to performance to security Selection= Differential weights*Benchmark's return = -0.1*2.9% - 0.1*1.6% + 0*0.5% = - 0.13%

Excess Performance = -0.36%-0.13% = -0.49%

F 221 .60% Actual Return Actual Weight Benchmark Weight Benchmark return 2 Equity 2.40% 0.6 0.7 2.90% 3 Bonds 1.40% 0.3 0.2 4B 2 Equity 3 Bonds 4 Cash Actual Return 0.024 0.014 0.005 Actual Weight 0.6 0.3 Benchmark Weight 0.7 0.2 Benchmark return 0.0